ANS: T PTS:4 The slope of a distance vs. time graph is the acceleration. ANS: A racecar driver steps on the gas, changing his speed from 10 m/s to 30 m/s in 4 seconds. The acceleration of the racecar is 10 m/s?. ANS: F When an object reaches its terminal speed, its acceleration is zero. ANS:T As you go higher and higher above the surface of the earth, the mass of a body stays constant. ANS: T As an artillery shell travels from the gun to the target, the projectiles vertical speed changes, but its horizontal speed stays constant. ANS: T 6. PTS: 4 7. PTS: 4 8. PTS: 4 9. PTS: 4 10. PTS: 4
Answer #1

6) False

slope of distance - time curve is dx/dt = speed of particle

7) False

acceleration is (a) = change in velocity/time

= (30-10)/4 =20/4 = 5m/s2

8) True

terminal velocity is constant velocity so acceleration is zero.

9) true

mass of a body is amount of matter of object. it is constant at everywhere in nature

(10) True

in horizental direction there is no force so speed of projectle is constant but in verticle direction due to gravity force velocity changes.